x + 10 5/8 = 16 1/4
Found 2 solutions by rfer, deborabr:
Answer by rfer(16322)

x=16 1/4-10 5/8
x=16 2/8-10 5/8
x=15 10/8-10 5/8
x=5 5/8

Answer by deborabr(181)

x + 10 5/8 = 16 1/4
x + 50/8 = 16/4
8x/8 + 50/8 = 32/8
8x + 50 = 32
8x = 32-50
8x = -18
x = -18/8
x = -9/4